# ATM: Adversarial Tuning Multi-agent System Makes a Robust Retrieval-Augmented Generator

### v2

@@ D.3 Infrastructure
We conduct SFT, DPO and MITO with full-parameter model optimization, which requires more GPU memory than our GPU can hold in its HBM with vanilla Distributed Data Parallel (DDP, Li et al., 2020). To this end, we use mixed-precision training with a model in bfloat16 data type Kalamkar et al. (2019) implemented in apex777https://github.com/NVIDIA/apex and ZeRO Rajbhandari et al. (2020) stage 1 implemented with DeepSpeed888https://github.com/microsoft/DeepSpeed which splits optimizer state across GPUs thus saving memory.
@@
We utilize vLLM999https://github.com/vllm-project/vllm with optimized Paged Attention for LLM inference Kwon et al. (2023) which is seamlessly compatible with the state-of-the-art LLM library transformers Wolf et al. (2020). For the fabrications generation, we selected the decoding hyper-parameter and top_p for fake knowledge fabrication in order to encourage generating diversity.
@@ Appendix E Affects of the Number of Documents
Appendix F Detailed Analysis of Attacker
In order to have an intuition of the increasing attack intensity of Attacker, we conduct case study and visualization to analyze its attacking patterns.
F.1 Visualization and Analysis
For the visualization analysis, we analyze the Log Loss of Generator given Attacker fabrications across different iterations. As depicted in Figure 6, the increasing PPL of the Generator shows that Attacker is emerging increasingly stronger attacking patterns which greatly obstruct the Generator from generating right answers.
F.2 Case Study
As shown in Table 7 with fabrications provided by the Attacker across different optimization iterations. Our provided question is: What character did Natalie Portman play in star wars? The proposed answer is: Padmé Amidala.

### v3

@@ D.3 Infrastructure
We conduct SFT, DPO and MITO with full-parameter model optimization, which requires more GPU memory than our GPU can hold in its HBM with vanilla Distributed Data Parallel (DDP, Li et al., 2020). To this end, we use mixed-precision training with a model in bfloat16 data type Kalamkar et al. (2019) implemented in apex666https://github.com/NVIDIA/apex and ZeRO Rajbhandari et al. (2020) stage 1 implemented with DeepSpeed777https://github.com/microsoft/DeepSpeed which splits optimizer state across GPUs thus saving memory.
@@
We utilize vLLM888https://github.com/vllm-project/vllm with optimized Paged Attention for LLM inference Kwon et al. (2023) which is seamlessly compatible with the state-of-the-art LLM library transformers Wolf et al. (2020). For the fabrications generation, we selected the decoding hyper-parameter and top_p for fake knowledge fabrication in order to encourage generating diversity.
@@ Appendix E Affects of the Number of Documents
Appendix F Case Study of Attacker
In order to have an intuition of the increasing attack intensity of Attacker, we conduct case study to analyze its attacking patterns. As shown in Table 7 with fabrications provided by the Attacker across different optimization iterations. Our provided question is: What character did Natalie Portman play in star wars? The proposed answer is: Padmé Amidala.